OpenStudy (anonymous):

If a = 3 & b = -4, find the value of 5ab - 3(a^2b)

OpenStudy (anonymous):

b=15ab ??

OpenStudy (jamesj):

If x = 5 and y = 7 what does x + y = ... ?

OpenStudy (anonymous):

so how do i solve to get what does x + y = ? do i need to multiply the two ? or divide ?

OpenStudy (jamesj):

Err .... x + y = 5 + 7 = 12

OpenStudy (jamesj):

Try another one. If x = 2 and y = -1, what is x + 2y^2 ?

OpenStudy (anonymous):

sounds like your getting real irratated, i'm sorry. >__< Grrrr' & it also sounds like your not wanting to just give me the answer, so thank you for that :]

OpenStudy (jamesj):

I want you to understand. Try this. Suppose x = 5. What is 2x + 1?

OpenStudy (jamesj):

Even easier then. Suppose x = 5. What is x + 1?

OpenStudy (jamesj):

Even easier than that even. What is 5 + 1?

OpenStudy (anonymous):

thank you so much! i want to understand this as well! 5 + 1 = 6

OpenStudy (jamesj):

Right. Now suppose x = 5. What is x + 1?

OpenStudy (anonymous):

7 ?

OpenStudy (jamesj):

If x = 5, then x + 1 = 5 + 1 = 6 yes?

OpenStudy (anonymous):

rite. i understand now. thank you:) so much!

OpenStudy (jamesj):

So if a = 3 & b = -4, what is 5ab ?

OpenStudy (anonymous):

ughhhh,,,,, >__<

OpenStudy (jamesj):

Let's go back a few steps again. Let x = 5. What is 2x + 1 ?

OpenStudy (jamesj):

Ok ... I'm out of here, sorry. Good luck.

OpenStudy (anonymous):

2x + 1 would be 26 because the 5 is the x , so put that with the 2 & plus it by the 1 makes it 26 ...? ok, thanks for all the help!